# Natural Language Generation from Visual Sequences: Challenges and Future Directions

### v1

@@ 4.2 The Vision-to-Language Bridge
Some V2L model architectures utilize an intermediate module that bridges the input and output modalities for effectively conditioning the text generation on the extracted visual features. Different models operationalize this module with different degrees of complexity. Earlier approaches for several multi-image-to-text tasks condition the text generation process by directly fusing vision encoder outputs with the language decoder input Kim et al. (2018). Some architectures employ cross-attention mechanisms to focus on the relevant parts of the visual features at various temporal positions during decoding Yao et al. (2015). However, approaches that adopt pre-trained models—e.g., CLIP-ViT-L Radford et al. (2021) as the visual model—tend to employ learnable intermediate layers for aligning and converting outputs of the vision encoder into a format that the language decoder can understand.
In some of the proposed models, this intermediate module is a single linear layer that transforms the visual features into a common shared space, which can be used by the language decoder Ko et al. (2023); Liu et al. (2023). In other models, advanced transformer-based projectors such as a Q-Former Li et al. (2023c) are used for their ability to leverage cross-modal interactions effectively Han et al. (2024). In essence, Q-Former uses dynamic query vectors that are pre-trained to attend to both visual and textual representations, enhancing its ability to generalize and perform well (relative to a single linear layer) across different tasks. Besides these popular methods for adapting multimodal information, some approaches make use of graph neural networks for capturing relationships between objects in the images at different temporal positions and words in the corresponding sentences of the text Zhang and Peng (2019). While there is no definitive way to design this intermediate module, recent work has compared the two approaches, i.e., using cross-attention between modalities or using a multimodal projector for transforming vision encoder features into the language space, and found that the latter leads to a stable/improved performance of models Laurençon et al. (2024b).
@@ 4.3 Language Decoder
Subsequent models have replaced RNNs with the transformer architecture owing to its computation scalability and efficiency in handling long context-windows. Besides the initial word embedding layer and the position encoding step (for maintaining information about the input sequence token order), a transformer decoder is typically made up of multiple identical blocks. Each block comprises a multi-head self-attention layer for modeling intra-sentence relations (between the words) and a multi-head cross-attention layer for learning relationships between representations of each word and the outputs of the visual encoder/projector. For instance, in the CC task, this refers to conditioning each word in the caption on vision encoder outputs (denoted as ‘difference-representations’).
@@
Instead of training the decoder from scratch, some approaches use language models such as GPT-2 Radford et al. (2019) and LLAMA 2 Touvron et al. (2023), which are pre-trained on several text-only tasks such as question-answering and text classification/completion. The pre-trained language models are either used directly for generation by freezing their parameters Han et al. (2023b, a, 2024), or by inserting and fine-tuning additional adaptive layers on top of them for ensuring relevance of the generated text to the downstream task of interest Yu et al. (2021). We also note that some models incorporate information from external knowledge bases/graphs into the decoder module to improve coherence and factuality of the generated text, e.g., TextKG Gu et al. (2023b) for the VC task and KG Story Hsu et al. (2020) for the VST task.
4.4 Off-the-shelf Pre-trained VLMs
The standard model architecture we have discussed so far is also present in more powerful general-purpose foundation VLMs (pre-trained on several tasks using large amounts of data), which can be used directly for multi-image-to-text tasks. Their pre-training process typically happens in two stages—self-supervised alignment training and visual instruction tuning. During the first stage, only the parameters of the intermediate module connecting both unimodal backbones are updated (commonly using paired image-text data) utilizing a contrastive training objective.
In the second stage, models are instruction-tuned using multi-turn conversations obtained for visual data either through crowd-sourcing or by leveraging tools such as GPT-4 OpenAI (2023). Contrary to task-specific modeling approaches, these pre-trained VLMs are simply prompted (typically in a zero-shot manner) using visual tokens accompanied by task-specific instructions. Some of the pre-trained VLMs that are used off-the-shelf for the multi-image-to-text tasks include: ViLA for VIDQA, mPLUG-2 for VC, VideoLLAMA for MAAD, and LLaVA-NeXT for VST Wang et al. (2025); Xu et al. (2023); Xie et al. (2024); Surikuchi et al. (2024).
5 Evaluation
Given all the similarities described above, it is not surprising that all multi-image-to-text tasks are also evaluated leveraging similar methods. These methods range from using traditional n-gram matching metrics to obtaining human judgments and ratings to, more recently, using off-the-shelf pre-trained VLMs assessing the generated output. We broadly classify these evaluation methods into two main categories—automatic and human evaluation. In the following subsections, we discuss the several quantitative metrics and benchmarks widely used for each of the tasks, along with the rationales for relying on them.
5.1 Automatic Evaluation
To computationally assess the quality of model-generated text along different aspects, several automatic metrics have been proposed. While some metrics rely on answers/text provided by human annotators, others are reference-free and assess model outputs independent of the ground-truth data. Besides computational metrics, the community has also relied on benchmark datasets designed to reveal various general capabilities of models.
Reference-based metrics.
The five multi-image-to-text tasks we examined primarily assessed model-generated candidate text by comparing it to corresponding human-written references. Specifically, traditional metrics that were originally designed for evaluating machine translation and text summarization tasks—BLEU Papineni et al. (2002), METEOR Banerjee and Lavie (2005), and ROUGE Lin (2004)—are used to measure precision and recall of overlapping $\mathit{n}$-grams between the candidate and reference text. Usually, metrics such as CIDEr Vedantam et al. (2015) and SPICE Anderson et al. (2016) that have been specifically developed for the evaluation of image and video captioning are also used in conjunction with the above three metrics.
All the above-mentioned metrics rely on direct raw text comparisons of ground-truth references and model outputs. As this ground-truth data might not be always available, embedding-level reference-based evaluation metrics such as WMD Kusner et al. (2015), BERTScore Zhang* et al. (2020), and ViLBERTScore Lee et al. (2020) have been proposed. Recent work for VC, VIDQA, and MAAD tasks have incorporated these metrics to measure the similarity between candidate and reference embeddings, obtained by projecting corresponding text into a common pre-trained semantic space Lin et al. (2022); Han et al. (2023a, 2024).
Reference-free metrics.
Comparing model-generated text to ground-truth references, typically provided by crowd-workers or scraped from the internet has various limitations. Most reference-based metrics do not account for the visual modality upon which the generated text is conditioned. Furthermore, as shown in Figure 2, text in the VC, MAAD, and VST tasks often complements the visual input by encompassing various abstract/creative concepts, and is not merely descriptive. This makes reference-based metrics generally inappropriate for accurately evaluating multi-image-to-text tasks.
For the reasons detailed above, various reference-free metrics such as MAUVE Pillutla et al. (2021) and UNION Guan and Huang (2020) have been proposed for unimodal open-ended text generation tasks. However, for the visually-conditioned text generation tasks, this adoption/shift is still relatively recent, with a persistent emphasis on reference-based $\mathit{n}$-gram metrics to date. That said, various reference-free metrics have been recently proposed to assess different aspects of evaluation that are important for several tasks. For instance, metrics such as CLIPScore Hessel et al. (2021) and GROOViST Surikuchi et al. (2023) have been developed for evaluating visual grounding—the degree of alignment between the generated text and the visual input—in VC and VST tasks. Similarly, the RoViST Wang et al. (2022) suite of metrics has been proposed to assess coherence, the extent of repetition, and visual grounding of the generated text. We note that there are also metrics such as CRITIC Han et al. (2024) and CM (character matching) Hong et al. (2023) designed to evaluate task-specific aspects such as the accuracy of referencing to characters in the model outputs. Besides the above-mentioned metrics (tailored to measure specific features of the generated content), there is also an increasing adoption and reliance on using pre-trained LLMs and VLMs as judges Li et al. (2024a). Essentially, these pre-trained general-purpose models are prompted to score or rate a model-generated response along any of the evaluation dimensions of interest, such as fluency or relevance (e.g., ‘How fluent is the generated text on a scale of 1 to 5?’). However, the effectiveness and reliability of this approach is still debated Bavaresco et al. (2024).
Benchmarks.
With the increase in scale, data, and extensive multi-step training processes it is difficult to fully understand the capabilities of models based only on their performance on held-out test splits of task-specific datasets. To address this limitation, numerous benchmark datasets have been proposed that focus on evaluating more fine-grained abilities of models. Some of the popular benchmarks proposed to test models trained for multi-image scenarios include: NLVR2 Suhr et al. (2019) which focuses on models’ ability to understand the visual compositionality given a pair of images and corresponding textual description; ViLMA Kesen et al. (2024) and MVBench Li et al. (2024c) which focus on testing models’ spatio-temporal reasoning capabilities (e.g., counting actions across frames of a video); Mementos Wang et al. (2024a) which studies object and behavioral hallucinations, and their interconnectedness.
Despite continued progress to improve and update existing benchmark datasets to cover various edge cases, evaluation using benchmarks is not without its limitations. For instance, some benchmarks are constructed using data from test/validation splits of existing popular datasets in the community, leading to a potential contamination problem Laurençon et al. (2024a). Moreover, recent modeling approaches typically tend to incorporate most existing benchmark datasets into their fine-tuning process to ensure the stability and generalization of models in real-world applications. When such approaches abstain from disclosing the data used for training the models, it generally undermines the process of automatic evaluation using benchmarks for comparing models against each other.
5.2 Human Evaluation
Given the current state of automatic evaluation, some multi-image-to-text tasks such as VC and VST rely on human evaluation to accurately determine the quality of the model-generated text. This process involves recruiting online crowd-workers who are native or proficient speakers of the target language. Depending on the type of data or variation of the task, annotators with expertise and familiarity with terminology relevant to the corresponding domain (e.g., medical or sports videos) might be preferred.
Participants of the evaluation study are provided with a set of task-specific rubrics/instructions along with representative examples required for judging the model outputs. They are asked to assess the overall quality of model-generated outputs either independently (per sample) Surikuchi et al. (2023) or relative to outputs from other models Wang et al. (2022). Alternatively, evaluators might be required to provide scores/ratings for various criteria ranging from broad (e.g., text conciseness, fluency, grammatical correctness) to specific (e.g., factuality, hallucinations, expressiveness). The obtained scores are usually compared pairwise to rank models appropriately.
Some pre-trained VLM frameworks such as LLaVA-RLHF Sun et al. (2024) leverage this qualitative feedback to optimize model parameters for learning to generate human-preferred text. Although human evaluation is still indispensable for several tasks, it is also expensive, time-consuming, and challenging. Defining clear evaluation protocols for ensuring the reliability and quality of human judgments is an active research area Kasai et al. (2022); Ruan et al. (2024).
6 Discussion
As discussed above, the problem of generating text from a sequence of temporally ordered images or frames is a challenging one, and relevant to several downstream tasks and applications. Here, we reflect on some crucial aspects and outline various prospective research directions (RDs) and takeaways.
RD 1: Towards more naturalistic scenarios.
Many of the multi-image-to-text tasks we consider in this work have real-world applications. For instance, solutions to the CC task can be used for assisted surveillance and for tracking changes in digital media assets Jhamtani and Berg-Kirkpatrick (2018). In the MAAD task, models are required to generate descriptions that complement information in the original audio dialog/soundtrack, for improved accessibility to visually impaired users and for enhancing the visual experience of sighted users Han et al. (2023a).
However, many day-to-day human-centered scenarios involve personalizing to various contexts or situations. We argue that existing multi-image-to-text tasks in their definitions and settings do not fully reflect this aspect. Tailoring model-generated descriptions/narrations to the perspective of end-users requires task settings in which models and humans can interact iteratively. Such settings would enable incorporation of human expectations and communicative contexts which typically tend to be dynamic in real-world applications. To this end, we advocate for variations of existing tasks where models can learn to contextualize and reason through interactions with other agents (humans or other models) for generating stories, descriptions, or answers. Furthermore, we also advocate for exploration of controlled task settings in which models are expected to generate text adhering to a specific style Yang and Jin (2023) or point-of-view.
RD 2: Are general-purpose VLMs all we need?
As discussed in Section 4.4, VLMs that are trained on various general-purpose datasets are increasingly being used for multi-image-to-text tasks through prompting. Powerful open-source models such as Molmo Deitke et al. (2024), and models optimized for multi-image scenarios such as Mantis Jiang et al. (2024) are becoming increasingly available, suggesting that the trend of adopting them off-the-shelf for solving many V2L tasks is widespread. General-purpose VLMs learn abundant information through multitask pre-training and have a modular design, making them suitable for many downstream tasks. Their modularity also enables for seamless adaptation of VLMs to various novel domains (e.g., medical science) by updating only a small fraction of their parameters Li et al. (2023a).
Despite the promising generalization of VLMs to certain tasks and domains, they have also been shown to be sensitive to prompts Liu et al. (2024); Schlarmann and Hein (2023) and biased towards the textual modality Rahmanzadehgervi et al. (2024). To address these problems, recent work proposes various prompt engineering techniques to facilitate inference-time adaptation of prompts to make them more suitable for the specific task of interest Ma et al. (2023); Gu et al. (2023a). On the other hand, task-specific model architectures consist of components designed to effectively address specialized aspects of the tasks, e.g., computing a difference representation of the input image pair in CC. We advocate for modular modeling approaches that bring together efficient task-specific components and combine them with the powerful foundational VLMs. Furthermore, we argue that using graph-based architectures and memory-based modules would result in improved tracking of entity positions/relationships and enable models to assign saliency to memorable events in tasks like VST or MAAD.
RD 3: Improving and rethinking evaluation.
In Section 5, we discussed the various approaches for evaluating model outputs in multi-image-to-text tasks. While human evaluation is impractical for conducting large-scale assessments, existing automatic evaluation metrics are limited in terms of fully capturing the abilities of models. Increasingly various benchmarking datasets are being proposed to assess models along different axes important for grounding language in the visual input Li et al. (2024b). However, many benchmarks often suffer from the problem of visual content irrelevance, which refers to models performing well on the benchmark datasets by primarily relying only on the language modality Chen et al. (2024a). Furthermore, data leakage and contamination problems (see section 5.1) also hinder fair and accurate testing of model’s skills using benchmarks.
While it is important to continue directing research efforts towards developing more extensive multi-image benchmarks such as ReMI Kazemi et al. (2024), we argue that the purpose of evaluation is to also provide insights that can be directly leveraged for improving model architectures and learning procedures. For single-image-to-text tasks, recent works have adapted interpretability methods that focus on understanding the behavior and internal representations of models Neo et al. (2024); Yu and Ananiadou (2024). These methods can complement traditional evaluation techniques for enabling intra- and inter-model comparisons of behaviors and mechanisms for obtaining a holistic understanding. We strongly advocate for work that adapts various categories of interpretability methods for multi-image-to-text scenarios.
7 Conclusion
In this position paper, we focused on the problem of multi-image V2L generation and connected the various tasks that are typically considered separate by the research community. We proposed a method for quantitatively exploring the current landscape of this problem, using which we uncovered relationships between task objectives and their corresponding datasets. Despite having different characteristics in terms of the objectives or the input-output data, we argued that all the tasks present a common set of challenges for developing models and for assessing their generated outputs. To understand the progress made over the years with regard to multi-image-to-text generation, we extensively reviewed the different modeling approaches and evaluation protocols pertaining to each of the tasks, and discussed them in a comprehensive and unified manner. As the problem of generating text conditioned on sequences of multiple temporally ordered images or video frames has various real-world applications, we underline the challenges and propose several concrete research directions informed by insights from linguistics, cognitive sciences, and natural language processing (NLP) aimed towards facilitating further advancements. We argue that leveraging these insights could also help the development of better VLMs, which are currently not immune from some of the highlighted limitations.
Acknowledgments
We are immensely grateful to our colleagues at the Dialogue Modelling Group (DMG) for their invaluable suggestions at different stages of this work. We thank Alberto Testoni and Anna Bavaresco for their insightful feedback. Raquel Fernández was supported by the European Research Council (ERC Consolidator Grant DREAM 819455).
Appendix A Quantitative Analysis of Tasks
For computing similarity scores, we used the image and text encoders of the ViT-B/32 version of the CLIP model. Obtained visual similarity and textual consistency scores are subtracted from $1.0$ for representing the tasks on the shared space in Figure 2. For the Multi-image/Video Question Answering task, we set textual consistency value to $1.0$ due to the unavailability of datasets having multiple-sentences in the textual outputs.
Appendix B Models
Table 3 outlines the models we reviewed for each of the multi-image-to-text tasks in this work. To reflect the evolution of models, we considered both RNN-based approaches and the more contemporary Transformer-based architectures that obtained state-of-the-art results.
{tblr}
colspec = X[c,h]X[c]X[c]X[c], stretch = 0, rowsep = 5pt, hlines = xkcdBlack, 0.5pt, vlines = xkcdBlack, 0.25pt
& Change Captioning, dataset: Spot-the-diff
Caption: The blue truck is no longer there. A car is approaching the parking lot from the right.
Video Question Answering, dataset: IntentQA Li et al. (2023b), Question: Why did the man point to the screen when talking to the child?
Answer: Draw child’s attention.
Video Captioning, dataset: Charades
Caption: The person is sitting at the dining room table wrapped in a blanket. The person is eating cereal and drinking orange juice.
Movie Auto Audio Description, dataset: MAD-v1
Story: Sully adjusts his seat harness. A male passenger looks up from his magazine … Sully sticks out an arm as the jet bellies down onto the river.
Visual Storytelling, dataset: VIST
Story: A discus got stuck up on the roof. Why not try getting it down with a soccer ball? … It didn’t work so we tried … are all stuck on the roof.
Table 2: Examples for each of the multi-image-to-text tasks we consider in this work.
@@ Appendix B Models
DUDA Park et al. (2019) ResNet, RNN NA RNN
Multi-image/Video Question Answering
@@
MM-Narrator Zhang et al. (2024) CLIP-ViT-L NA GPT-4

### v2

@@ 2.2 The Vision-to-Language Bridge
Some vision-to-language model architectures use an intermediate module that bridges the input and output modalities for effectively conditioning text generation on the extracted visual features. Different models operationalize this module with different degrees of complexity. Earlier approaches condition the text generation process by directly fusing vision encoder outputs with the language decoder input Kim et al. (2018). Some architectures employ cross-attention mechanisms to focus on the relevant parts of the visual features at various temporal positions during decoding Yao et al. (2015). However, approaches that adopt pre-trained models—e.g., CLIP-ViT-L Radford et al. (2021) as the visual model—tend to employ learnable intermediate layers for aligning and converting outputs of the vision encoder into a format that the language decoder can understand.
In some of the proposed models, this intermediate module is a single linear layer that transforms the visual features into a common shared space, which can then be used by the language decoder Ko et al. (2023); Liu et al. (2023). In other models, advanced transformer-based projectors such as a Q-Former Li et al. (2023b) are used based on their ability to leverage cross-modal interactions effectively Han et al. (2024). In essence, Q-Former uses dynamic query vectors that are pre-trained to attend to both visual and textual representations, enhancing its ability to generalize and perform well (relative to a single linear layer) across different tasks. Besides these popular methods for adapting multimodal information, some approaches make use of graph neural networks for capturing relationships between objects in the images at different temporal positions and words in the corresponding sentences of the text Zhang and Peng (2019).
While there is no definitive way to design this intermediate module, recent work has compared the two approaches, i.e., using cross-attention between modalities or using a multimodal projector for transforming vision encoder features into the language space, and found that the latter leads to improvements in performance of models Laurençon et al. (2024). This suggests that using intermediate multimodal projectors can help in extracting meaningful visual representations.
@@ 2.3 Language Decoder
Subsequent models have replaced RNNs with the transformer architecture owing to its computational scalability and efficiency in handling long context-windows. Besides the initial word embedding layer and the position encoding step (for maintaining information about the input sequence token order), a transformer decoder is typically made up of multiple identical blocks. Each block comprises a multi-head self-attention layer for modeling intra-sentence relations (between the words) and a multi-head cross-attention layer for learning relationships between representations of each word and the outputs of the visual encoder or projector. For instance, in the Change Captioning task, this refers to conditioning each word in the caption on vision encoder outputs (denoted as ‘difference-representations’).
@@
Instead of training the decoder from scratch, some approaches use language models such as GPT-2 Radford et al. (2019) and LLAMA 2 Touvron et al. (2023), which are pre-trained on several text-only tasks such as question-answering and text classification. The pre-trained language models are either used directly for generation by freezing their parameters Han et al. (2023b, a, 2024), or by inserting and fine-tuning additional adaptive layers on top of them for ensuring relevance of the generated text to the downstream task of interest Yu et al. (2021). We also note that some models incorporate information from external knowledge bases into the decoder module to improve coherence and factuality of the generated text, e.g., TextKG Gu et al. (2023b) for the Video Captioning task and KG Story Hsu et al. (2020) for the Visual Storytelling task. This approach is shown to improve models in terms of correctness of answers, relevance of descriptions, and expressiveness of generated narratives. On the other hand, methods relying on external knowledge face several practical limitations, such as retrieval latency and model scalability. Furthermore, for applications in some domains (e.g., medical, financial), interacting with external sources may not be feasible due to privacy or security constraints.
@@
DUDA ResNet, RNN NA RNN
Video QA
@@
MM-Narrator Zhang et al. (2024a) CLIP-ViT-L NA GPT-4